# RIs-Calib: An Open-Source Spatiotemporal Calibrator for Multiple 3D Radars and IMUs Based on Continuous-Time Estimation

### v1

@@ Abstract
Aided inertial navigation system (INS), typically consisting of an inertial measurement unit (IMU) and an exteroceptive sensor, has been widely accepted as a feasible solution for navigation. Compared with vision-aided and LiDAR-aided INS, radar-aided INS could achieve better performance in adverse weather conditions since the radar utilizes low-frequency measuring signals with less attenuation effect in atmospheric gases and rain. For such a radar-aided INS, accurate spatiotemporal transformation is a fundamental prerequisite to achieving optimal information fusion. In this work, we present RIs-Calib: a spatiotemporal calibrator for multiple 3D radars and IMUs based on continuous-time estimation, which enables accurate spatiotemporal calibration and does not require any additional artificial infrastructure or prior knowledge. Our approach starts with a rigorous and robust procedure for state initialization, followed by batch optimizations, where all parameters can be refined to global optimal states steadily. We validate and evaluate RIs-Calib on both simulated and real-world experiments, and the results demonstrate that RIs-Calib is capable of accurate and consistent calibration. We open-source our implementations at (https://github.com/Unsigned-Long/RIs-Calib) to benefit the research community.
@@ I Introduction and Related Works
Inertial navigation systems (INSs), utilizing inertial measurement units (IMUs) as sensing modality, can provide six-degrees-of-freedom motion estimation in the three-dimensional space. However, long-term drift typically exists in INSs due to the noise and biases in inertial measurements. A feasible solution to combat this issue is integrating exteroceptive sensors, such as a camera or light detection and ranging (LiDAR), into INSs, i.e., constructing aided INSs. While vision-aided INSs [1] or LiDAR-aided INSs [2] could achieve accurate ego-motion estimation, they are highly vulnerable to adverse weather, such as fog, rain, and snow. Conversely, radar-aided INSs [3, 4, 5] are insensitive to such challenging conditions as radars utilize lower-frequency signals which have lighter attenuation effect in atmospheric gases and rain. Due to this fact, radar-aided INSs have attracted significant research efforts in recent years. For such systems, accurate spatiotemporal calibration is highly required since ill-calibrated spatiotemporal parameters could significantly affect the fusion performance.
@@
In contrast to radar-related calibration, sufficient motion excitation is required in IMU-related calibration when collecting data to guarantee parameter observability. Benefiting from such dynamic calibration, it allows temporal determination. To calibrate the extrinsic, intrinsic, and temporal parameters of a Camera/IMU suite, Hu et al. [13] employed a multi-state constrained invariant Kalman filter to perform online calibration, which greatly reduces the computational complexity compared with the traditional Kalman filter. Similarly, Huai et al. [14] proposed a keyframe-based visual-inertial odometry with online self-calibration, where a detailed observability analysis for spatiotemporal parameters is also performed. It should be noted that the above methods are all discrete-time-based ones, where trajectories are represented by a series of discrete poses in estimation, and simplifying assumptions are generally needed, which would introduce ineluctable errors in calibration.
@@
Different from the discrete-time trajectory representation, continuous-time representation employs continuous-time functions to encode trajectories, where poses can be computed at any arbitrary time, making it well-suited for asynchronous or high-frequency sensor fusion. The most representative work is the well-known Kalibr [15], which first employs B-splines as the continuous-time trajectory representation to perform spatiotemporal calibration for a global shutter camera and an IMU. Subsequently, building upon Kalibr, the rolling shutter camera is further supported in [16] by Huai et al. Similarly employing B-splines as the continuous-time trajectory representation, Lv et al. [17] proposed a targetless LiDAR/IMU calibration method. In their further work [18], observability-aware modules are leveraged to address degenerate motions.
@@
While 3D radar-inertial navigation systems [3, 4, 5] have attracted significant research interest and have been increasingly developed recently, gaps exist in accurate spatiotemporal calibration for such sensor suites. To this end, based on continuous-time optimization, we propose a targetless spatiotemporal calibrator for sensor suites that integrates multiple 3D radars and IMUs. Specifically, we perform a rigorous initialization procedure to obtain initial guesses of states in the estimator, which requires no prior knowledge. Subsequently, based on the initialized parameters and raw measurements from radars and IMUs, we form a nonlinear factor graph including radar factors and IMU factors, and optimize it for several batches until the final convergence. The main contribution of this work can be summarized as follows:
@@
    Different from traditional continuous-time-based calibration methods that employ continuous-time curves to represent trajectories, we innovatively employ them to encode rotation and velocity curves, which is naturally compatible with the measurements of radar and IMU, and effectively reduces optimization complexity.
@@
    We carried out both simulated and real-world experiments to demonstrate the high accuracy and repeatability of the proposed method.
  4. 4.
    We open-source our implementations to benefit the community. To the best of our knowledge, this is the first open-source multi-radar multi-IMU spatiotemporal calibration toolbox.
@@ II-A Notation
For a sensor suite that integrates $\mathcal{N}_{b}$ IMUs and $\mathcal{N}_{r}$ radars, we consider $\{b^{i}\}$ and $\{r^{j}\}$ as frames of the $i$-th IMU and the $j$-th radar respectively, whose measurement sequences are denoted as $\mathcal{D}(b^{i})$ and $\mathcal{D}(r^{j})$, where $i\in\left[0,\cdots,\mathcal{N}_{b}\right)$ and $j\in\left[0,\cdots,\mathcal{N}_{r}\right)$. We employ the Euclidean parametrization to represent the 3D rigid body transformation from frame $\{a\}$ to frame $\{b\}$ as follows:
@@
where ${{}^{b}_{a}\boldsymbol{T}}\in\mathrm{SE(3)}$ is also known as the transform matrix. ${{}^{b}_{a}\dot{\boldsymbol{R}}}\Leftrightarrow{{}^{b}_{a}\boldsymbol{\omega}}$, ${{}^{b}_{a}\ddot{\boldsymbol{R}}}\Leftrightarrow{{}^{b}_{a}\dot{\boldsymbol{\omega}}}$, ${{}^{b}\dot{\boldsymbol{p}}_{a}}\Leftrightarrow{{}^{b}\boldsymbol{v}_{a}}$, and ${{}^{b}\ddot{\boldsymbol{p}}_{a}}\Leftrightarrow{{}^{b}\dot{\boldsymbol{v}}_{a}}$ are the angular velocity, angular acceleration, linear velocity, and linear acceleration of frame $\{a\}$ with respect to and parameterized in frame $\{b\}$, all of which live in $\mathbb{R}^{3}$. Finally, we denote $\hat{(\cdot)}$ as the noisy measurement, $\boldsymbol{\Sigma}_{(\cdot)}$ as the corresponding covariance matrix of a residual, and $\rho_{(\cdot)}$ as the Cauchy loss function to reduce the influence of outliers in least-squares problems.
@@ II-B Sensor Model
${{}^{g^{i}}}\hat{\boldsymbol{\omega}}(\tau)$ $={{}^{g^{i}}_{b^{i}}\boldsymbol{R}}\cdot{{}^{b^{i}}}\boldsymbol{\omega}(\tau)+\boldsymbol{b}_{\omega}^{i}+\boldsymbol{\varepsilon}_{\omega}^{i}=\boldsymbol{f}_{\omega}\left({{{}^{b^{i}}}\boldsymbol{\omega}}(\tau),\boldsymbol{x}_{(b^{i},in)}\right)$ (2)
@@
$\boldsymbol{x}_{\left(b^{i},in\right)}=\left\{\boldsymbol{b}_{\omega}^{i},\boldsymbol{b}_{a}^{i},{{}^{g^{i}}_{b^{i}}\boldsymbol{R}}\right\}$ (3)
where $\{g^{i}\}$ and $\{a^{i}\}$ are the sensor frames of the gyroscope and accelerometer, respectively; ${{{}^{b^{i}}}\boldsymbol{\omega}}(\tau)$ and ${{{}^{b^{i}}}\boldsymbol{a}}(\tau)$ are the ideal angular velocity and linear acceleration in frame $\{b^{i}\}$; $\boldsymbol{b}_{\omega}^{i}$ and $\boldsymbol{b}_{a}^{i}$ are the gyroscope and accelerometer biases, which are modeled as random walks and can be considered constant when the calibration data is short; $\boldsymbol{\varepsilon}_{\omega}^{i}$ and $\boldsymbol{\varepsilon}_{a}^{i}$ denote the measurement noises; ${{}^{g^{i}}_{b^{i}}\boldsymbol{R}}$ is the rotational misalignment between the accelerometer and gyroscope by assuming that frame $\{b^{i}\}$ coincides with frame $\{a^{i}\}$.
@@ II-C Continuous-time Representation
To accurately fuse asynchronous and high-frequency measurements from multiple radars and IMUs, we employ the uniform B-splines as the continuous-time representation to encode the velocity and rotation trajectories. Compared to other continuous-time functions, e.g., Gaussian process regression and hierarchical wavelets, the uniform B-splines have closed-from analytic derivatives and local controllability, which could reduce numerical errors and yield a sparse estimator in optimization.
Specifically, given a series of temporally uniformly distributed velocity control points $\boldsymbol{v}_{i}$, $\boldsymbol{v}_{i+1}$, $\cdots$, $\boldsymbol{v}_{i+d}$, the velocity B-spline of degree $d$ could be expressed as:
@@
As for the rotation B-spline, its control points which live in Lie Group $\mathrm{SO(3)}$ should be mapped to Lie Algebra $\mathfrak{so}(3)$ for scalar multiplication. Concretely, the rotation B-spline of degree $d$ could be written as:
@@ III-A Problem Formulation
The structure of the proposed spatiotemporal calibration method for multiple radars and IMUs is shown in Fig. 2. The system starts with a rigorous initialization procedure, which recovers the rotation and velocity B-splines of a so-called virtual central IMU (i.e., the reference IMU whose frame is denoted as $\{b^{c}\}$ in this work), as well as the extrinsics and gravity. Subsequently, a nonlinear factor graph that minimizes radar residuals, IMU residuals, and center residuals would be formed and optimized several times until the final convergence.
The full state vector $\boldsymbol{x}$ in the system includes extrinsics $\boldsymbol{x}_{\left(\;\cdot\;,\;ex\right)}$, temporal parameters $\boldsymbol{x}_{\left(\;\cdot\;,\;tm\right)}$, intrinsics $\boldsymbol{x}_{\left(\;\cdot\;,\;in\right)}$, a set of control points in the rotation and velocity B-splines of the central IMU $\boldsymbol{x}_{\left(cp,rot\right)}$ and $\boldsymbol{x}_{\left(cp,vel\right)}$, and the gravity vector ${}^{b^{c}_{0}}\boldsymbol{g}$ parameterized in the first frame of the central IMU $\{b^{c}_{0}\}$, which is defined as:
@@
$\boldsymbol{x}_{(cp,rot)}$ =$=\left\{\cdots,{{}^{b^{c}_{0}}_{b^{c}_{k}}\boldsymbol{R}},\cdots\right\}$ $\boldsymbol{x}_{(cp,vel)}$ =$=\left\{\cdots,{{}^{b^{c}_{0}}\boldsymbol{v}_{b^{c}_{k}}},\cdots\right\}$
where ${{}^{b^{c}}_{b^{i}}}\boldsymbol{R}$, ${{}^{b^{c}}}\boldsymbol{p}_{b^{i}}$, and ${{}^{b^{c}}\tau_{b^{i}}}$ are the extrinsic rotation, extrinsic translation, and time offset of the $i$-th IMU with respect to the central IMU, respectively; ${{}^{b^{c}}_{r^{j}}}\boldsymbol{R}$, ${{}^{b^{c}}}\boldsymbol{p}_{r^{j}}$, and ${{}^{b^{c}}\tau_{r^{j}}}$ are the extrinsic rotation, extrinsic translation, and time offset of the $j$-th radar with respect to the central IMU, respectively; ${{}^{b^{c}_{0}}_{b^{c}_{k}}\boldsymbol{R}}$ and ${{}^{b^{c}_{0}}\boldsymbol{v}_{b^{c}_{k}}}$ are the $k$-th control points in the rotation and velocity B-splines respectively with $k\in\left[0,\cdots,\mathcal{N}_{cp}\right)$; ${{}^{b^{c}_{0}}\boldsymbol{g}}\in\mathbb{R}^{3}$ is the two-degrees-of-freedom gravity vector with a constant magnitude $||{{}^{b^{c}_{0}}\boldsymbol{g}}||\approx 9.81m/s^{2}$. Note that we parameterize all spatiotemporal parameters with respect to the virtual central IMU $\{b^{c}\}$, and consider $\{b^{c}_{0}\}$ as the reference frame of the B-splines, i.e., the static world frame.

### v2

@@ Abstract
Aided inertial navigation system (INS), typically consisting of an inertial measurement unit (IMU) and an exteroceptive sensor, has been widely accepted as a feasible solution for navigation. Compared with vision-aided and LiDAR-aided INS, radar-aided INS could achieve better performance in adverse weather conditions since the radar utilizes low-frequency measuring signals with less attenuation effect in atmospheric gases and rain. For such a radar-aided INS, accurate spatiotemporal transformation is a fundamental prerequisite to achieving optimal information fusion. In this work, we present RIs-Calib: a spatiotemporal calibrator for multiple 3D radars and IMUs based on continuous-time estimation, which enables accurate spatiotemporal calibration and does not require any additional artificial infrastructure or prior knowledge. Our approach starts with a rigorous initialization procedure to recover spatiotemporal parameters and kinematic B-splines of the sensor suite from the raw measurements. Following that, several batch optimizations would be conducted, where all parameters would be refined to global optimal states steadily. We validate and evaluate RIs-Calib on both simulated and real-world experiments, and the results demonstrate that RIs-Calib is capable of accurate and consistent calibration. We open-source our implementations at (https://github.com/Unsigned-Long/RIs-Calib) to benefit the research community.
@@ I Introduction and Related Works
Inertial navigation systems (INSs), utilizing inertial measurement units (IMUs) as sensing modality, can provide six-degrees-of-freedom motion estimation in the three-dimensional space. However, long-term drift typically exists in INSs due to the noise and biases in inertial measurements. A feasible solution to combat this issue is integrating exteroceptive sensors, such as a camera or light detection and ranging (LiDAR), into INSs, i.e., constructing aided INSs. While vision-aided INSs [1, 2] or LiDAR-aided INSs [3, 4] could achieve accurate ego-motion estimation, they are highly vulnerable to adverse weather, such as fog, rain, and snow. Conversely, radar-aided INSs [5, 6, 7] are insensitive to such challenging conditions as radars utilize lower-frequency signals which have lighter attenuation effect in atmospheric gases and rain [8]. Due to this fact, radar-aided INSs have attracted significant research efforts in recent years. For such systems, accurate spatiotemporal calibration is highly required since ill-calibrated spatiotemporal parameters could significantly affect the fusion performance [9].
@@
In contrast to radar-related calibration, sufficient motion excitation is required in IMU-related calibration when collecting data to guarantee parameter observability. Benefiting from such dynamic calibration, it allows temporal determination. To calibrate the extrinsic, intrinsic, and temporal parameters of a Camera/IMU suite, Hu et al. [17] employed a multi-state constrained invariant Kalman filter to perform online calibration, which greatly reduces the computational complexity compared with the traditional Kalman filter. Similarly, Huai et al. [18] proposed a keyframe-based visual-inertial odometry with online self-calibration, where a detailed observability analysis for spatiotemporal parameters is also performed. The above filter-based methods are discrete-time-based ones, which represents trajectories using discrete poses in estimator, thus simplified kinematic assumptions are generally needed when processing asynchronous measurements, which would introduce ineluctable errors in calibration.
@@
Different from the discrete-time trajectory representation, continuous-time representation employs continuous-time functions to encode trajectories, where poses can be computed at any arbitrary time, making it well-suited for asynchronous or high-frequency sensor fusion. The most representative work is the well-known Kalibr [19], which first employs B-splines as the continuous-time trajectory representation to perform spatiotemporal calibration for an IMU and a global shutter camera. Subsequently, building upon Kalibr, the rolling shutter camera is further supported in [20] by Huai et al. Similarly employing B-splines as the continuous-time trajectory representation, Lv et al. [21] proposed a targetless LiDAR/IMU calibration method. In their further work [22], observability-aware modules are leveraged to address degenerate motions.
The continuous-time estimation had also been employed in our previous work [9, 23] for targetless camera/IMU and LiDAR/IMU spatiotemporal calibration. Different from them, the work presented in this paper focuses on 3D millimeter-wave radar/IMU spatiotemporal calibration, and faces several challenges compared with camera/IMU and LiDAR/IMU calibration: the structurelessness, sparsity, and significant noise of radar target measurements. This leads to difficulties in data association and continuous-time trajectory recovery if we still adhere to the pipeline of our previous work [9, 23]. Considering this, we cleverly and reasonably abandon the previous practice of employing continuous-time representation to model the ego-motion trajectory (rotation and position curves), and instead employ it to model rotation and ego-velocity kinematic curves. Such modeling can rigorously and conveniently fuse both inertial measurements from IMUs and Doppler velocity measurements from radars in the estimator for spatiotemporal calibration, while avoiding challenging data association for radar measurements [24, 25] and potential inconsistency and even association errors.
@@
In summary, while 3D radar-inertial navigation systems [5, 6, 7] have attracted significant research interest and have been increasingly developed recently, gaps exist in accurate spatiotemporal calibration for such sensor suites. To this end, based on continuous-time optimization, we propose a targetless spatiotemporal calibrator for sensor suites that integrates multiple 3D radars and IMUs. Specifically, we perform a rigorous initialization procedure to obtain initial guesses of states in the estimator, which requires no prior knowledge. Subsequently, based on the initialized parameters and raw measurements from radars and IMUs, we form a nonlinear factor graph including radar factors and IMU factors, and optimize it for several batches until the final convergence. The main contributions in our work can be listed as follows:
@@
    Different from traditional continuous-time-based calibration methods that employ B-splines to represent time-varying pose (rotation and position), we innovatively employ them to encode rotation and velocity curves, which is naturally compatible with the measurements of both radar and IMU, and effectively reduces optimization complexity.
@@
    We carried out both simulated and real-world experiments to demonstrate the high accuracy and repeatability of the proposed method. We open-source our implementations to benefit the community.
@@ II-A Notation
For a sensor suite that integrates $\mathcal{N}_{b}$ IMUs and $\mathcal{N}_{r}$ radars, we consider $\{b^{i}\}$ and $\{r^{j}\}$ as frames of the $i$-th IMU and the $j$-th radar respectively. The measurements of $\{b^{i}\}$ and $\{r^{j}\}$ are denoted as $\mathcal{D}(b^{i})$ and $\mathcal{D}(r^{j})$, where $i\in\left[0,\cdots,\mathcal{N}_{b}\right)$ and $j\in\left[0,\cdots,\mathcal{N}_{r}\right)$. We employ the Euclidean parametrization to represent the 3D transformation from frame $\{a\}$ to coordinate frame $\{b\}$ as follows:
@@
where ${{}^{b}_{a}\boldsymbol{T}}\in\mathrm{SE(3)}$ is the transform matrix. ${{}^{b}_{a}\dot{\boldsymbol{R}}}\Leftrightarrow{{}^{b}_{a}\boldsymbol{\omega}}$, ${{}^{b}_{a}\ddot{\boldsymbol{R}}}\Leftrightarrow{{}^{b}_{a}\dot{\boldsymbol{\omega}}}$, ${{}^{b}\dot{\boldsymbol{p}}_{a}}\Leftrightarrow{{}^{b}\boldsymbol{v}_{a}}$, and ${{}^{b}\ddot{\boldsymbol{p}}_{a}}\Leftrightarrow{{}^{b}\dot{\boldsymbol{v}}_{a}}$ are the angular velocity, angular acceleration, linear velocity, and linear acceleration of frame $\{a\}$ with respect to and parameterized in frame $\{b\}$, all of which live in $\mathbb{R}^{3}$. Finally, we denote $\hat{(\cdot)}$ as the noisy measurement, $\boldsymbol{\Sigma}_{(\cdot)}$ as the corresponding covariance matrix of a residual, and $\rho_{(\cdot)}$ as the Cauchy loss function to reduce the influence of outliers in least-squares problems.
@@ II-B Sensor Model
${{}^{g^{i}}}\hat{\boldsymbol{\omega}}(\tau)$ $={{}^{b^{i}}}\boldsymbol{\omega}(\tau)+\boldsymbol{b}_{\omega}^{i}+\boldsymbol{\varepsilon}_{\omega}^{i}=\boldsymbol{f}_{\omega}\left({{{}^{b^{i}}}\boldsymbol{\omega}}(\tau),\boldsymbol{x}_{(b^{i},in)}\right)$ (2)
@@
$\boldsymbol{x}_{\left(b^{i},in\right)}=\left\{\boldsymbol{b}_{\omega}^{i},\boldsymbol{b}_{a}^{i}\right\}$ (3)
where $\{g^{i}\}$ and $\{a^{i}\}$ are the sensor frames of the gyroscope and accelerometer, respectively; ${{{}^{b^{i}}}\boldsymbol{\omega}}(\tau)$ and ${{{}^{b^{i}}}\boldsymbol{a}}(\tau)$ are the ideal angular velocity and linear acceleration in frame $\{b^{i}\}$; $\boldsymbol{b}_{\omega}^{i}$ and $\boldsymbol{b}_{a}^{i}$ are the gyroscope and accelerometer biases, which are modeled as random walks and can be considered constant when the calibration data is short; $\boldsymbol{\varepsilon}_{\omega}^{i}$ and $\boldsymbol{\varepsilon}_{a}^{i}$ denote the measurement noises.
@@ II-C Continuous-time Representation
To accurately fuse asynchronous and high-frequency measurements from multiple radars and IMUs, we employ the uniform B-splines as the continuous-time representation to encode the velocity and rotation trajectories. Compared to other continuous-time functions, e.g., Gaussian process [26], hierarchical wavelets [27], and Chebyshev interpolation [28], the uniform B-splines have closed-from analytic derivatives and local controllability, which could yield a sparse system and reduce computational complexity in optimization.
Specifically, given a sequence of velocity control points $\boldsymbol{v}_{i}$, $\boldsymbol{v}_{i+1}$, $\cdots$, $\boldsymbol{v}_{i+d}$ that are temporally uniformly distributed, the velocity B-spline of degree $d$ could be expressed as:
@@
As for the rotation B-spline, its control points which live in Lie Group $\mathrm{SO(3)}$ should be mapped to Lie Algebra $\mathfrak{so}(3)$ for scalar multiplication. Concretely, the $d$-degree rotation B-spline can be expressed as:
@@ III-A Problem Formulation
The structure of the proposed spatiotemporal calibration method for multiple radars and IMUs is shown in Fig. 2. The system starts with a rigorous initialization procedure, which recovers extrinsics, world-frame gravity vector, as well as the velocity and rotation B-splines of a so-called virtual central IMU (namely the reference IMU, it’s denoted as $\{b^{c}\}$). Subsequently, a nonlinear factor graph that minimizes IMU residuals and radar residuals would be optimized several batches until the final convergence.
The full state vector $\boldsymbol{x}$ in the system includes extrinsics $\boldsymbol{x}_{\left(\;\cdot\;,\;ex\right)}$, time offsets $\boldsymbol{x}_{\left(\;\cdot\;,\;tm\right)}$, intrinsics $\boldsymbol{x}_{\left(\;\cdot\;,\;in\right)}$, a set of control points in the rotation and velocity B-splines of the central IMU $\boldsymbol{x}_{\left(cp,rot\right)}$ and $\boldsymbol{x}_{\left(cp,vel\right)}$, and the gravity vector ${}^{b^{c}_{0}}\boldsymbol{g}$ parameterized in the first frame of the central IMU $\{b^{c}_{0}\}$, which is defined as:
@@
$\boldsymbol{x}_{(cp,rot)}$ =$=\left\{\cdots,{{}^{b^{c}_{0}}\boldsymbol{R}_{k}},\cdots\right\}$ $\boldsymbol{x}_{(cp,vel)}$ =$=\left\{\cdots,{{}^{b^{c}_{0}}\boldsymbol{v}_{k}},\cdots\right\}$
where ${{}^{b^{c}}_{b^{i}}}\boldsymbol{R}$, ${{}^{b^{c}}}\boldsymbol{p}_{b^{i}}$, and ${{}^{b^{c}}\tau_{b^{i}}}$ are the extrinsic rotation, extrinsic translation, and time offset of the $i$-th IMU with respect to the central IMU, respectively; ${{}^{b^{c}}_{r^{j}}}\boldsymbol{R}$, ${{}^{b^{c}}}\boldsymbol{p}_{r^{j}}$, and ${{}^{b^{c}}\tau_{r^{j}}}$ are the extrinsic rotation, extrinsic translation, and time offset of the $j$-th radar, respectively; The first frame of the reference IMU, i.e., coordinate frame $\{b^{c}_{0}\}$, is considered as the static world frame; ${{}^{b^{c}_{0}}\boldsymbol{R}_{k}}$ and ${{}^{b^{c}_{0}}\boldsymbol{v}_{k}}$ are the $k$-th control points of the rotation and velocity B-splines respectively, with $k\in\left[0,\cdots,\mathcal{N}_{cp}\right)$; ${{}^{b^{c}_{0}}\boldsymbol{g}}\in\mathbb{R}^{3}$ is the world-frame two-degrees-of-freedom gravity vector with a constant magnitude $||{{}^{b^{c}_{0}}\boldsymbol{g}}||\approx 9.81m/s^{2}$. Note that all spatiotemporal parameters are with respect to the virtual central IMU $\{b^{c}\}$, while other time-varying kinematic states are expressed with respect to $\{b^{c}_{0}\}$.